# On the realization of infinite (continuous)

### v2

@@ 6 Summary
In this paper we have studied the algebraic aspects of the Lagrangian formulation for the infinite spin field model proposed in [12]. The formulation is constructed in the framework of the BRST approach to higher spin field theory. The main elements of this approach are the Fock space vectors $|{\Psi}\rangle$ (2.1) and the operator constraints (2.5)-(2.8) in this space, forming a closed algebra in terms of commutators, which ensures gauge invariance of the Lagrangian theory. The operator constraints in the case under consideration includes the radius $R$ of $\rm{AdS}_{4}$ space and the arbitrary real massive parameter $\boldsymbol{\mu}$ associated with the continuous representation. It was proved that the above operator constraints are the invariants of the $\mathrm{SO}(2,3)$ group and the Casimir operators of the $\mathfrak{so}(2,3)$ algebra in the Fock space are completely expressed through these constraints in the form (5.24) and (5.36).
@@ Acknowledgments
Authors are grateful to R.R. Metsaev for stimulating discussions and comments. The authors are also thankful to M.A. Vasiliev and Yu.M. Zinoviev for useful comments. M.A.P. is grateful to A.A. Golubtsova for helpful discussions. The work of M.A.P. was supported by the JINR AYSS Foundation, Project No 24-301-03.

### v3

@@ 6 Summary
In this paper we have studied the algebraic aspects of the Lagrangian formulation for the infinite spin field model proposed in [12]. The main elements of our analysis are the fields ${\Phi}$ (2.1) and the operator constraints (2.5)-(2.8) acting on these fields and forming a closed algebra in terms of commutators, which ensures gauge invariance of the Lagrangian theory. The operator constraints in the case under consideration includes the radius $R$ of $\rm{AdS}_{4}$ space and the arbitrary real massive parameter $\boldsymbol{\mu}$ associated with the continuous representation. It was proved that the above operator constraints are the invariants of the $\mathrm{SO}(2,3)$ group and the Casimir operators of the $\mathfrak{so}(2,3)$ algebra in the space of vectors (2.1) are completely expressed through these constraints in the form (5.24) and (5.36).
@@ Acknowledgments
Authors are grateful to R.R. Metsaev for stimulating discussions and comments. The authors are also thankful to M.A. Vasiliev, Yu.M. Zinoviev and V.A. Krykhtin for useful comments. M.A.P. is grateful to A.A. Golubtsova for helpful discussions. The work of M.A.P. was supported by the JINR AYSS Foundation, Project No 24-301-03.